OpenStudy (kaitlyn_nicole):

Write the equation of the line that passes through (3, –2) and has a slope of 4 in point-slope form.

RhondaSommer (rhondasommer):

\[y=mx+b\] m is slope so (4)

RhondaSommer (rhondasommer):

then pug in your (x,y) coordinate to get (b)

RhondaSommer (rhondasommer):

and you have yourself an equation :)

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ % (a,b) &({\color{red}{ 3}}\quad ,&{\color{blue}{ -2}})\quad \end{array} \\\quad \\ % slope = m slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies4 \\ \quad \\ % point-slope intercept y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values}\\ \qquad \uparrow\\ \textit{point-slope form}\) plug them in be happy eat ice-cream do the hula hoop

rebeccaxhawaii (rebeccaxhawaii):

plug it into slope form

OpenStudy (kaitlyn_nicole):

what do you mean plug in to get b?

OpenStudy (kaitlyn_nicole):

like solve it, right?

RhondaSommer (rhondasommer):

so \[y=4x+b\]\[x=3\]\[y=-2\]\[-2=4(3)+b\]

RhondaSommer (rhondasommer):

so solve that

OpenStudy (jdoe0001):

nothing to solve really you're given "m", the slope, of 4 you're also given "x" and "y" and you don't need to change the form

OpenStudy (kaitlyn_nicole):

-2=12+b -14=b?

RhondaSommer (rhondasommer):

yep :)

RhondaSommer (rhondasommer):

\[y=4x+(b)

OpenStudy (kaitlyn_nicole):

thank you!

RhondaSommer (rhondasommer):

yeah :)

OpenStudy (jdoe0001):

yw

OpenStudy (kaitlyn_nicole):

wait, it doesnt say that on my answers though? maybe we over simplified it? y + 2 = 4(x – 3) y – 3 = 4(x + 2) x – 3 = 4(y + 2) x + 2 = 4(y – 3)

OpenStudy (kaitlyn_nicole):

so would it be D?

RhondaSommer (rhondasommer):

um; i dont think we over simplified

RhondaSommer (rhondasommer):

ohhh

RhondaSommer (rhondasommer):

\[y-y^1=m(x-x^1)\]

OpenStudy (kaitlyn_nicole):

so b?

RhondaSommer (rhondasommer):

\[y^1=-2\]\[x^1=3\] and \[m~still=4\]

RhondaSommer (rhondasommer):

so \[y-(-2)or(y+2)=4(x-3)\]

OpenStudy (kaitlyn_nicole):

oh a

RhondaSommer (rhondasommer):

that make sense? i was using a different form XD

RhondaSommer (rhondasommer):

yep :)